OpenStudy (anonymous):

A sample of a gas occupies a volume of 752 mL at 25°C. What volume will the gas occupies if the temperature increases to 50°C, if the pressure remains constant?

OpenStudy (cuanchi):

V1/T1=V2/T2 isolate V2 and calculate the temperature has to be in Kelvin K=C+273